On average, 3 accidents happen every week along Great North Road. What is the probability that we will observe 9 accidents in a month? 


1
Expert's answer
2021-11-23T18:00:56-0500

Let "x=" the number of accidents happened: "x\sim Po(\lambda t)"

Given "\lambda=3, t=4"


"P(x=9)=\dfrac{e^{-3(4)}\cdot(3(4) )^9}{9!}=0.087364"

The probability that we will observe 9 accidents in a month is 0.087364.